OpenStudy (anonymous):

The population of New Haven was 148,220 in 2005. Since then the population has been decreasing at an annual rate of 2.9%. If this rate on decline continues, what will the population be in the year 2015? Y=a*b^x where a is the starting amount; b is the decay rate (decreasing subtract from 1); x is time.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

\(\sf Y=a*b^x\) a = 148,220 b = .029 x = 10

OpenStudy (anonymous):

Thank you so much!! Could you help me with another one?? :) @sleepyjess

OpenStudy (sleepyjess):

Yep :)

OpenStudy (anonymous):

Suppose the population of deer in a region was 3500 in the year 2000. Since then the population has grown by annually. What will the approximate population be in the year 2020? Y=a*b^x where a is starting amount; b is the growth rate (increasing add to 1); x is the time. @sleepyjess

OpenStudy (sleepyjess):

Has grown by what annually?

OpenStudy (sleepyjess):

I know that a = 2500 and x = 20

OpenStudy (sleepyjess):

WAIT! On the first one, the decay rate should be subtracted from 1, so b = .971

OpenStudy (anonymous):

Haha okay. And I dont know what grows annually or whatever. :/ That is the question im stuck on :(

OpenStudy (sleepyjess):

It doesn't give a value?

OpenStudy (anonymous):

What I posted is what is on the paper.

OpenStudy (sleepyjess):

hmmmm

OpenStudy (anonymous):

I will try to take a picture and post it. One sec.

OpenStudy (sleepyjess):

ok, I have to go for about 25-30 minutes, but while I'm gone @AnswerMyQuestions or @paki may be able to help.

OpenStudy (anonymous):

Here is the question. Please excuse my bad handwriting hah.

OpenStudy (anonymous):

Not bad compared to mine... >.>

OpenStudy (anonymous):

Haha well could you help me with the question? Pleasee

OpenStudy (anonymous):

@answermyquestions

OpenStudy (anonymous):

I am not sure. ;C

OpenStudy (anonymous):

Man :((

OpenStudy (anonymous):

@paki could you help me?

OpenStudy (anonymous):

He's offline. :(

OpenStudy (anonymous):

@Directrix or @Nurali might help you.

OpenStudy (anonymous):

Aww :'(

OpenStudy (anonymous):

I sure hope so